Boat A leaves the pier at 9 a.m. at 8 Knots. A half hour later, Boat B leaves the same pier in the same direction traveling at 10 knots. At what time will Boat B overtake Boat A?
Answer by stanbon(75887)

Boat A leaves the pier at 9 a.m. at 8 Knots. A half hour later, Boat B leaves the same pier in the same direction traveling at 10 knots. At what time will Boat B overtake Boat A?
---------
Boat A DATA::
rate = 8 kt/hr ; time = x hrs ; distance = 8x kts
--------
Boat B DATA:
rate = 10 kt/hr ; time = x-1/2 hrs ; distance = 10(x-(1/2)) = 10x-5
---------
Equation:
distance = distance
8x = 10x-5
2x = 5
x = 2.5 hrs
-------
Your Problem:
9:00 + 2.5 hrs = 11:30
=============================
Cheers,
Stan H.
===================